OpenStudy (anonymous):

Please someone explain step by step Q1:area: find area of a right angled triangle with side lengths 27cm, 35cm and 44.2cm. Q2:a can contains 375ml of soft drink and must have a height of 120mm. what will the diameter need to be? Q3:a garden hose has an internal diameter of 15 mm and a length of 5m. what volume of water fills the hose?

OpenStudy (wolf1728):

Q1 multiply the 2 "legs" of the triangle hypotenuse = 44.2 there fore the 2 legs are 27 and 35 27*35 = 945 cm²

OpenStudy (jdoe0001):



OpenStudy (wolf1728):

UGH !!! forgot to divide by two !!

OpenStudy (jdoe0001):

 \(\bf \textit{volume of a cylinder}=\pi r^2 h\quad volume=375\quad h=120\quad r=radius=\frac{diameter}{2} \\ \quad \\ 375=\pi r^2\cdot 120\) solve for "r" to get the diameter, keep in mind that radius is just half of the diameter